OpenStudy (solomonzelman):

IMPORTANT, PLEASE MAJOR HELP NEEDED!!! TRIGONOMETRY RVW. PLEASE IF YOU KNOW THIS STAFF!

OpenStudy (solomonzelman):

@kiergurl5, there you go! @myininaya @Zarkon @robtobey @timo86m IMPORTANT!!!!

OpenStudy (anonymous):

Thanks so much, @SolomonZelman! :)

OpenStudy (solomonzelman):

@kiergurl5, ask questions about specific concepts, I;ll find and msg other helpers.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

Do you know the basics about the circles parabolas and hyperbolas?

OpenStudy (anonymous):

kind of... not completly

OpenStudy (solomonzelman):

lets start with circles.\[(x-h)^2+(y-k)^2=r^2\]\[(h,k)~~~~~~~~is~~the~~center.\]\[r~~~~~~~~is~~~the~~radius.\]

OpenStudy (solomonzelman):

can you write the equation for this circle?

OpenStudy (anonymous):

ummm.....

OpenStudy (solomonzelman):

(I won't always be available, But I want to make sure the user gets help needed) @kiergurl5, try you can do it, compare to the formula.

OpenStudy (anonymous):

(0,0) is the center... so...\[(3-0)^{2}+(3-0)^{2}=r^2\]

OpenStudy (anonymous):

maybe?

OpenStudy (anonymous):

ok... 18=r^2

OpenStudy (anonymous):

\[\sqrt{18}=r\]

OpenStudy (solomonzelman):

Hold on, I mis-understood what you wrote, y and x remain the same,

OpenStudy (solomonzelman):

\[(x-0)^2+(y-0)^2=3^2\]\[(x)^2+(y)^2=3^2\]\[x^2+y^2=3^2\]\[x^2+y^2=9\]

OpenStudy (solomonzelman):

I am not substituting 3 for x and y, I am substituting 3 for r.

OpenStudy (solomonzelman):

UNDERSTAND?

OpenStudy (anonymous):

so i just did it backwards?

OpenStudy (solomonzelman):

You plugged in 3 for x and y,. but you are/were supposed to lug in 3 for r and (0,0) for (h,k)

OpenStudy (anonymous):

Ohhh....

OpenStudy (solomonzelman):

Lets do more examples.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

A little trickier....